# Multi-User Personalisation in Human-Robot Interaction: Using Quantitative Bipolar Argumentation Frameworks for Preferences Conflict Resolution

### v1

@@ Abstract
While personalisation in Human-Robot Interaction (HRI) has advanced significantly, most existing approaches focus on single-user adaptation, overlooking scenarios involving multiple stakeholders with potentially conflicting preferences. To address this, we propose the Multi-User Preferences Quantitative Bipolar Argumentation Framework (MUP-QBAF), a novel multi-user personalisation framework based on Quantitative Bipolar Argumentation Frameworks (QBAFs) that explicitly models and resolves multi-user preference conflicts. Unlike prior work in Argumentation Frameworks, which typically assumes static inputs, our approach is tailored to robotics: it incorporates both users’ arguments and the robot’s dynamic observations of the environment, allowing the system to adapt over time and respond to changing contexts. Preferences, both positive and negative, are represented as arguments whose strength is recalculated iteratively based on new information. The framework’s properties and capabilities are presented and validated through a realistic case study, where an assistive robot mediates between the conflicting preferences of a caregiver and a care recipient during a frailty assessment task. This evaluation further includes a sensitivity analysis of argument base scores, demonstrating how preference outcomes can be shaped by user input and contextual observations. By offering a transparent, structured, and context-sensitive approach to resolving competing user preferences, this work advances the field of multi-user HRI. It provides a principled alternative to data-driven methods, enabling robots to navigate conflicts in real-world environments.
keywords:
Assistive Robots , Multi-user Personalisation , Argumentation Frameworks , Gradual Argumentation , Conflict Resolution
[…]
[iri]organization=Institut de Robòtica i Informàtica Industrial (CSIC-UPC), addressline=Llorens i Artigas 4-6, city=Barcelona, postcode=08028, country=Spain
[…]
[iiia]organization=Institut d’Investigació en Intel·ligència Artificial (IIIA-CSIC), addressline=Campus de la UAB, city=Cerdanyola del Vallès, postcode=08193, country=Spain
1 Introduction
@@
This work introduces the Multi-User Preferences Quantitative Bipolar Argumentation Frameworks (MUP-QBAFs), a novel framework that enables robots to resolve conflicts when multiple users express competing preferences. The proposed method leverages Quantitative Bipolar Argumentation Frameworks (QBAFs) to select a preference by evaluating users’ stated preferences, their underlying reasons, and dynamic environmental observations, all modelled as arguments within the framework. The selected preference then parameterises the robot’s decision-making system. This framework offers additional advantages for HRI. Firstly, a key property of AFs is their representation as a graph, where the arguments are the nodes and the edges are their relations, producing transparent decisions [Rago_kr20], unlike opaque learning-based methods. Transparency in the decision-making process is fundamental to provide explanations to users, which is crucial in real-world HRI. Secondly, AFs are inherently adaptable; arguments can be added or modified dynamically, enabling immediate decision updates without retraining, making them suitable for long-term interactions where preferences evolve over time [Irfan_UMUAI23]. Finally, while most of the literature in AFs assumes static decisions, HRI introduces a unique requirement: iterative decision-making, that is, making the same decision several times during an interaction. The outcome of the decision can change from one iteration to another due to the actions of the robot, new observations, or new arguments provided by the users [Ferretti_AI17]. These dynamic properties are defined and demonstrated through a case study, in which an assistive robot must reconcile conflicting preferences between a caregiver and a care recipient during an assistive task. Finally, the possible outcomes of the framework are analysed, complementing it with the study of the arguments’ base score. The overall system architecture is illustrated in Fig. 1.
@@ 2 Preliminaries
2.1 Introduction of Argumentation Frameworks
@@
With the attack relation between arguments, the concept of defence appeared. For example, in Fig. 2(a), A attacks B and B attacks C, thus A is defending C since it attacks an argument that attacks C. With those definitions, the semantics for accepting arguments were introduced, being a set of arguments $S\subseteq\mathcal{A}$:
@@
The supporting arguments brought new semantics and definitions of relations, such as direct/indirect attackers or supporters of an argument, the safety sets, which are the ones that do not attack and support the same argument, and new admissible and preferred extensions. In the example from Fig. 2(b), the set {A, C} is not safe since the set is attacking and supporting B.
(a) AF with A attacking B and B attacking C.
(b) AF with A attacking B and C supporting B.
Figure 2: Argumentation Framework Examples.
2.2 Quantitative Bipolar Argumentation Frameworks
@@ 3 Related work
4 Multiple users’ preference selection system
This section introduces the modules (see Fig. 1) of our framework used to select a preference in a scenario with multiple users with possibly competing preferences for a specific action. Firstly, the robot observations and their history generate the task and users’ arguments in the arguments module (see Sec. 4.1). Next, these arguments are fed into the preference selector module formalised as a QBAF to determine the selected preference (see Sec. 4.2). Finally, the selected preference is input to the action sequence executor module to parameterise the action. It is important to note that developing a decision-making system to select the best action in a given state is outside the scope of this work.
@@ 4 Multiple users’ preference selection system
4.1 Robot observations and arguments modules
@@
4.2 Preference selector module
This module serves as the core computational engine for resolving conflicts and determining the robot’s course of action. Upon receiving the users’ preferences and arguments, the module initiates a four-stage process. Firstly, it analyses the preferences to detect any conflicts (see Sec. 4.2.1). Second, it formalises the conflict using our proposed AF (see Sec. 4.2.2). Third, a dedicated algorithm processes this framework to select a winning preference and decide the parameters of the robot’s action (see Sec. 4.2.3). Finally, we demonstrate that this decision-making process adheres to several properties that ensure its responsiveness in dynamic contexts (see Sec. 4.2.4).
4.2.1 Preferences and conflict types
@@
    $\exists o\in O.\ \exists U^{\prime}\subseteq U.\ |U^{\prime}|\geq 1\land\forall i\in U^{\prime}.\ (o,+)\in\hat{p}(i)\land\forall o^{\prime}\in O.\ o\neq o^{\prime}\to$
    $(o^{\prime},+)\notin\hat{p}(i)\land\big(\forall j\in U\setminus U^{\prime}.\ (o,0)\in\hat{p}(j)\land(o^{\prime},+)\notin\hat{p}(j)\big)$
@@
    $\begin{split}\exists\ i,j\in U.\ \exists\ o,o^{\prime}\in O.\ o^{\prime}\neq o\ \land\ i\neq j\ \land\ (o,+)\in\hat{p}(i)\ \land\ (o^{\prime},+)\in\hat{p}(j)\\ \land\ (o,+)\notin\hat{p}(j)\ \land\ (o^{\prime},+)\notin\hat{p}(i)\end{split}$
@@
    $\begin{split}\exists\ i,j\in U.\ \exists\ o,o^{\prime}\in O.\ o^{\prime}\neq o\ \land\ i\neq j\ \land\ (o,-)\in\hat{p}(i)\ \land\ (o^{\prime},-)\in\hat{p}(j)\\ \land\ (o,-)\notin\hat{p}(j)\ \land\ (o^{\prime},-)\notin\hat{p}(i)\end{split}$
@@
4.2.2 The MUP-QBAF framework
The arguments received from the arguments module, and defined in the previous section (see Sec. 4.1), are used to construct a Multi-User Preferences Quantitative Bipolar Argumentation Framework (MUP-QBAF):
@@ 4.2.2 The MUP-QBAF framework
4.2.3 Multi-user preference selector algorithm
@@
The first step of the algorithm is to calculate the final strengths of the MUP-QBAF’s arguments (line 1). The next step consists of checking if any of the conditions for conflicting scenarios from Sec. 4.2.1 are satisfied (line 2). In case of conflict, the algorithm will select an option $o^{*}\in O^{\prime}$ by running $g$ (line 3), which is a function that selects among the set of options $O^{\prime}\subseteq O$ with the highest strength: $o^{*}=g\{o|\sigma(o)=\max_{o^{\prime}\in O}\sigma(o^{\prime}),\mathcal{M},\nu\}$. Those options with the maximum strength are input into the $g$ function, which selects an option given some criteria. If there is only one option in $g$, the conflict is considered resolved, and the selected option is that one.
@@
4.2.4 Properties of MUP-QBAFs dynamic adaptation
@@
5 Using MUP-QBAF in robotic frailty assessment
@@
Firstly, the use case, the role of the robot, and the possible conflicts are explained (see Sec. 5.1). Then, an example of a possible AF (see Sec. 5.2) is provided. Moreover, some examples related to adaptation to dynamic context, which is crucial in HRI, are presented (see Sec. 5.3). Finally, an analysis of the proposed AF in different scenarios, and the sensitivity of the arguments’ base score is shown (see Sec. 5.4).
5.1 Clinical scenario and motivation
@@
5.2 Constructing the Argumentation Framework
@@
5.3 Dynamic adaptation examples
@@
5.4 Framework evaluation and sensitivity analysis
@@
Table 1: Distribution of preference selections for Repeat (R), No Repeat (N-R) and Tie across different argument configurations for the two conflicting scenarios. The first and fourth rows correspond to all the possible argument combinations. The second and fifth rows are the combinations where the argument T1 is activated. The third and last rows are the combinations when the risk of falling arguments (the upper part of the AF) are removed. All the arguments’ base score is 0.5. R means Repeat, and N-R means No Repeat.
Table 1 presents the selection percentages for various combinations of activated and deactivated arguments across both scenarios. The first row lists all possible combinations, the second row focuses on instances when the risk of falling argument (T1) is activated, and the third row shows the distributions when the risk of falling arguments are not included. In this last scenario, while the argument CR1 remains in the framework, it does not influence the final strengths because there is no path connecting it to the Option Arguments. From this table, we can observe the impact of the risk of falling argument (T1), which skews the distributions towards “no repeat” thereby making the assessment safer based on actual observations of the robot. This effect is particularly noticeable when comparing the values between the first and second rows for each scenario. In contrast, the rows that exclude the risk of falling arguments show a more balanced distribution between repeating and not repeating. This balance is expected since the Option Arguments have an equal number of attackers and supporters in both scenarios. The differences observed in the last row regarding repetition are attributable to the indirect attackers and supporters of the Option Arguments.
Additionally, it is possible to compute the influence of the relations between arguments on the Option Arguments, based on the approach from [Yin_ijcai24] named Relation Attribution Explanations (RAEs). This analysis shows the influence of each relation on the Option Arguments. Table 2 shows an example of influences for scenario (2), since it contains more arguments and is more complete, with all the arguments activated and base scores at 0.5.
@@
Table 2: Scenario 2 (caregiver for repeat, care recipient against repeat). Approximate influence of each relation between arguments on the final Option Arguments when all arguments are active and with a base score of 0.5. The value 0 is for strict zeros, and 0.00 refers to values <5<5 · $10^{-3}$.
@@
(a) Mean and standard deviation (95%) distribution of the difference between the final strengths of repeat and no repeat, ($\sigma(r)-\sigma(\neg r)$), for all the combinations in scenario (1). The top plot corresponds to the totality of the combinations (activated and deactivated), the second plot corresponds to the combinations where T1 is activated, and the third one corresponds to the combinations where T1 is deactivated.
(b) Percentages of the decisions, in scenario (1), where the caregiver is against repeat. The distribution of all combinations is plotted at the top. At the bot, the combinations where the risk of falling argument is active (T1). The blue dashed section refers to the percentages that the robot decided to repeat the tests. The yellow dashed section refers to not repeating. The green area refers to ties.
@@
Finally, we conducted a sensitivity analysis to evaluate how the base score of a critical argument influences the final decision. We varied the base score of the “risk of falling” argument (T1), which acts as a parameter to make the robot’s decision-making more or less conservative. Fig. 6(a) illustrates the results, showing the difference in final strength between the repeat and no repeat options ($\sigma(r)-\sigma(\neg r)$) as the base score of T1 is increased from 0 to 1 (all other arguments were held at a base score of 0.5). The plot shows a clear trend: as the T1 base score increases, the strength difference in favour of repeating diminishes. This directly impacts the robot’s choices, as shown in Fig. 6(b). The selection percentage for repeat decreases as the T1 base score grows. This confirms that a higher base score for the “risk of falling” argument makes the robot more cautious, eventually reaching a point where it will strictly select the no repeat option when this risk is present.
6 Limitations and outlook
@@
Regarding the AF design, the following limitations are noted: firstly, this work does not implement Argument Mining, which is the field in CA that generates and relates arguments from user input. The arguments in the AF presented in this work were decided through co-design with healthcare professionals. However, in a complete and fully autonomous real-world deployment, Argument Mining will be necessary for getting the users and task arguments. This will become an important challenge for future work, which can be potentially overcome with the usage of LLM [Gorur_arxiv24]. Secondly, all the base scores in Sec. 5 are set to 0.5, except in the sensitivity analysis. In some scenarios, it may seem reasonable for the system to assign different base scores to users’ arguments. For instance, users might give arguments that are considered very relevant and important to them [Battaglia_icsum24]. Those arguments could have a greater base score. Another factor considered when setting the base scores can be the expertise of the users in the context, e.g. in assistive scenarios, health-related arguments given by the caregiver may be more grounded than those given by the care recipient. While this can improve decision quality, it also raises a risk of systemic bias, prioritising caregiver preferences at the expense of the care recipient’s autonomy. Thirdly, while this study assumes distinct arguments for each user, scenarios where users share arguments but assign different base scores require further exploration. Methods to aggregate individual AFs and base scores into a general AF should be further explored [Dickie_AAMAS25].
Respecting the system’s interaction limitations, we mention the following: firstly, no explanation is generated regarding the reason for the robot’s selection. Explanations could be based on the influence of each argument or their relationships to the final decision (see Sec. 5.4). Similarly, if a user complains or questions the reason behind the robot’s decision, the robot must provide a clear and understandable explanation to facilitate the necessary feedback for adjusting its decision. To achieve this, it is essential to identify the necessary and sufficient changes and explanations required within the framework  [Kampik_IJAR24]. Secondly, this work is presented in a particular assistive use case. Even though it is generalisable for any scenario where multiple users have competing preferences, it should be tested and studied in different use cases and in a real-world scenario. Additionally, this work is based on a single-interaction approach, in which a care recipient interacts with the robot once, only to perform the tests. More investigation on how this system behaves in long-term interactions is needed.

### v2

# Multi-User Personalisation in Human-Robot Interaction: Resolving Preference Conflicts Using Gradual Argumentation

@@ Abstract
While personalisation in Human-Robot Interaction (HRI) has advanced significantly, most existing approaches focus on single-user adaptation, overlooking scenarios involving multiple stakeholders with potentially conflicting preferences. To address this, we propose the Multi-User Preferences Quantitative Bipolar Argumentation Framework (MUP-QBAF), a novel multi-user personalisation framework that extends Quantitative Bipolar Argumentation Frameworks (QBAFs) to model and resolve preference conflicts. Unlike prior work in Argumentation Frameworks, which typically assumes static inputs, our approach incorporates the users’ arguments and the robot’s dynamic observations of the environment, allowing the system to continuously adapt and respond to changing contexts. Preferences are represented as arguments whose strength is recalculated iteratively based on new information. The framework’s properties and capabilities are presented and validated through a realistic case study, where an assistive robot mediates between the conflicting preferences of a caregiver and a care recipient during a frailty assessment task. This evaluation includes a sensitivity analysis of argument base scores, demonstrating how preference outcomes can be shaped by user input and contextual observations. By offering a transparent, structured, and context-sensitive approach to resolving competing user preferences, this work advances the field of multi-user HRI, providing a principled alternative to data-driven methods, enabling robots to navigate conflicts in real-world environments.
I Introduction
@@
This work introduces the Multi-User Preferences Quantitative Bipolar Argumentation Frameworks (MUP-QBAFs), a novel framework that enables robots to resolve conflicts when multiple users express competing preferences. The proposed method extends Quantitative Bipolar Argumentation Frameworks (QBAFs) to select a preference by evaluating users’ stated preferences, their underlying reasons, and dynamic environmental observations, all modelled as arguments within the framework. This gradual argumentation approach enables a possible fine-tuning of users’ importance of each argument, and avoids the current limitation in abstract argumentation, which is the loss of information due to the framework reductions [Battaglia_icsum24]. The selected preference then parameterises the robot’s decision-making system. This framework offers additional advantages for HRI. Firstly, a key property of AFs is their representation as a graph, where the arguments are the nodes and the edges are their relations, producing transparent decisions [Rago_kr20], unlike opaque learning-based methods. Transparency in the decision-making process is fundamental to providing explanations to users, which is crucial in real-world HRI. Secondly, AFs are inherently adaptable; arguments can be added or modified dynamically, enabling immediate decision updates without retraining, making them suitable for long-term interactions where preferences evolve over time [Irfan_UMUAI23]. Finally, while most of the literature in AFs assumes static decisions, HRI introduces a unique requirement: iterative decision-making, that is, making the same decision several times during an interaction. The outcome of the decision can change from one iteration to another due to the actions of the robot, new observations, or new arguments provided by the users [Ferretti_AI17]. These dynamic properties are defined and demonstrated through a case study, in which an assistive robot must reconcile conflicting preferences between a caregiver and a care recipient during an assistive task. Finally, the possible outcomes of the framework are analysed, complementing it with the study of the arguments’ base score. The overall system architecture is illustrated in Fig. 1.
@@ II Preliminaries
II-A Introduction of Argumentation Frameworks
@@
With the attack relation between arguments, the concept of defence appeared. For example, in Fig. 2a, A attacks B and B attacks C, thus A is defending C since it attacks an argument that attacks C. With those definitions, the semantics for accepting arguments were introduced, being a set of arguments $S\subseteq\mathcal{A}$:
@@
The supporting arguments brought new semantics and definitions of relations, such as direct/indirect attackers or supporters of an argument, the safety sets, which are the ones that do not attack and support the same argument, and new admissible and preferred extensions. In the example from Fig. 2b, the set {A, C} is not safe since the set is attacking and supporting B.
(a) AF with only attacks.
(b) A bipolar AF.
Figure 2: Argumentation Framework Examples. In a), A is attacking B and B is attacking C; therefore, A is defending C. In b), A is attacking B, and C is supporting B.
II-B Quantitative Bipolar Argumentation Frameworks
@@ III Related work
IV Multiple users’ preference selection system
This section introduces the modules (see Fig. 1) of our framework used to select a preference in a scenario with multiple users with possibly competing preferences for a specific action. Firstly, the robot observations and their history generate the task and users’ arguments in the arguments module (see Sec. IV-A). Next, these arguments are fed into the preference selector module formalised as a QBAF to determine the selected preference (see Sec. IV-B). Finally, the selected preference is input to the action sequence executor module to parameterise the action. It is important to note that developing a decision-making system to select the best action in a given state is outside the scope of this work.
@@ IV Multiple users’ preference selection system
IV-A Robot observations and arguments modules
@@
IV-B Preference selector module
This module serves as the core computational engine for resolving conflicts and determining the robot’s course of action. Upon receiving the users’ preferences and arguments, the module initiates a four-stage process. Firstly, it analyses the preferences to detect any conflicts (see Sec. IV-B1). Second, it formalises the conflict using our proposed AF (see Sec. IV-B2). Third, a dedicated algorithm processes this framework to select a winning preference and decide the parameters of the robot’s action (see Sec. IV-B3). Finally, we demonstrate that this decision-making process adheres to several properties that ensure its responsiveness in dynamic contexts (see Sec. IV-B4).
IV-B1 Preferences and conflict types
@@
    $\exists o\in O.\;\exists U^{\prime}\subseteq U.\;|U^{\prime}|\geq 1\land\forall i\in U^{\prime}.\;(o,+)\in\hat{p}(i)$
    $\land\forall o^{\prime}\in O.\;o\neq o^{\prime}\to(o^{\prime},+)\notin\hat{p}(i)$
    $\land\big(\forall j\in U\setminus U^{\prime}.\;(o,0)\in\hat{p}(j)\land(o^{\prime},+)\notin\hat{p}(j)\big)$
@@
    $\exists i,j\in U.\;\exists o,o^{\prime}\in O.\;o^{\prime}\neq o\land i\neq j$
    $\land(o,+)\in\hat{p}(i)\land(o^{\prime},+)\in\hat{p}(j)$
    $\land(o,+)\notin\hat{p}(j)\land(o^{\prime},+)\notin\hat{p}(i)$
@@
    $\exists i,j\in U.\ \exists\ o,o^{\prime}\in O.\ o^{\prime}\neq o\land i\neq j$
    $\land\ (o,-)\in\hat{p}(i)\ \land\ (o^{\prime},-)\in\hat{p}(j)$
    $\land\ (o,-)\notin\hat{p}(j)\ \land\ (o^{\prime},-)\notin\hat{p}(i)$
@@
IV-B2 The MUP-QBAF framework
The arguments received from the arguments module, and defined in the previous section (see Sec. IV-A), are used to construct a Multi-User Preferences Quantitative Bipolar Argumentation Framework (MUP-QBAF):
@@ IV-B2 The MUP-QBAF framework
IV-B3 Multi-user preference selector algorithm
@@
The first step of the algorithm is to calculate the final strengths of the MUP-QBAF’s arguments (line 1). The next step consists of checking if any of the conditions for conflicting scenarios from Sec. IV-B1 are satisfied (line 2). In case of conflict, the algorithm will select an option $o^{*}\in O^{\prime}$ by running $g$ (line 3), which is a function that selects among the set of options $O^{\prime}\subseteq O$ with the highest strength: $o^{*}=g\{o|\sigma(o)=\max_{o^{\prime}\in O}\sigma(o^{\prime}),\mathcal{M},\nu\}$. Those options with the maximum strength are input into the $g$ function, which selects an option given some criteria. If there is only one option in $g$, the conflict is considered resolved, and the selected option is that one.
@@
IV-B4 Properties of MUP-QBAFs dynamic adaptation
@@
V Using MUP-QBAF in robotic frailty assessment
@@
Firstly, the use case, the role of the robot, and the possible conflicts are explained (see Sec. V-A). Then, an example of a possible AF (see Sec. V-B) is provided. Moreover, some examples related to adaptation to dynamic context, which is crucial in HRI, are presented (see Sec. V-C). Finally, an analysis of the proposed AF in different scenarios, and the sensitivity of the arguments’ base score is shown (see Sec. V-D).
V-A Clinical scenario and motivation
@@
V-B Constructing the Argumentation Framework
@@
V-C Dynamic adaptation examples
@@
V-D Framework evaluation and sensitivity analysis
@@
TABLE I: Distribution of preference selections for Repeat (R), No Repeat (N-R) and Tie across different argument configurations for the two conflicting scenarios. The first and fourth rows correspond to all the possible argument combinations. The second and fifth rows are the combinations where the argument T1 is activated. The third and last rows are the combinations when the risk of falling arguments (the upper part of the AF) are removed. All the arguments’ base score is 0.5. R means Repeat, and N-R means No Repeat.
Table I presents the selection percentages for various combinations of activated and deactivated arguments across both scenarios. The first row lists all possible combinations, the second row focuses on instances when the risk of falling argument (T1) is activated, and the third row shows the distributions when the risk of falling arguments are not included. In this last scenario, while the argument CR1 remains in the framework, it does not influence the final strengths because there is no path connecting it to the Option Arguments. From this table, we can observe the impact of the risk of falling argument (T1), which skews the distributions towards “no repeat” thereby making the assessment safer based on actual observations of the robot. This effect is particularly noticeable when comparing the values between the first and second rows for each scenario. In contrast, the rows that exclude the risk of falling arguments show a more balanced distribution between repeating and not repeating. This balance is expected since the Option Arguments have an equal number of attackers and supporters in both scenarios. The differences observed in the last row regarding repetition are attributable to the indirect attackers and supporters of the Option Arguments.
Additionally, it is possible to compute the influence of the relations between arguments on the Option Arguments, based on the approach from [Yin_ijcai24] named Relation Attribution Explanations (RAEs). This analysis shows the influence of each relation on the Option Arguments. Table II shows an example of influences for scenario (2), since it contains more arguments and is more complete, with all the arguments activated and base scores at 0.5.
@@
TABLE II: Scenario 2 (caregiver for repeat, care recipient against repeat). Approximate influence of each relation between arguments on the final Option Arguments when all arguments are active and with a base score of 0.5. The value 0 is for strict zeros, and 0.00 refers to values <5<5 · $10^{-3}$.
@@
(a) Mean and standard deviation (95%) distribution of the difference between the final strengths of repeat and no repeat, ($\sigma(r)-\sigma(\neg r)$), for all the combinations in scenario (1). The top plot corresponds to all combinations (activated and deactivated), the second plot corresponds to combinations where T1 is activated, and the third one to combinations where T1 is deactivated.
(b) Percentages of the decisions, in scenario (1), where the caregiver is against repeat. The top plot shows all combinations. At the bottom, combinations where the risk of falling argument (T1) is active. The blue dashed section refers to percentages where the robot decided to repeat the tests; the yellow dashed section refers to not repeating; the green area refers to ties.
@@
Finally, we conducted a sensitivity analysis to evaluate how the base score of a critical argument influences the final decision. We varied the base score of the “risk of falling” argument (T1), which acts as a parameter to make the robot’s decision-making more or less conservative. Fig. 6a illustrates the results, showing the difference in final strength between the repeat and no repeat options ($\sigma(r)-\sigma(\neg r)$) as the base score of T1 is increased from 0 to 1 (all other arguments were held at a base score of 0.5). The plot shows a clear trend: as the T1 base score increases, the strength difference in favour of repeating diminishes. This directly impacts the robot’s choices, as shown in Fig. 6b. The selection percentage for repeat decreases as the T1 base score grows. This confirms that a higher base score for the “risk of falling” argument makes the robot more cautious, eventually reaching a point where it will strictly select the no repeat option when this risk is present.
VI Limitations and outlook
@@
Regarding the AF design, the following limitations are noted: firstly, this work does not implement Argument Mining, which is the field in CA that generates and relates arguments from user input. The arguments in the AF presented in this work were decided through co-design with healthcare professionals. However, in a complete and fully autonomous real-world deployment, Argument Mining will be necessary for getting the users and task arguments. This will become an important challenge for future work, which can be potentially overcome with the usage of LLM [Gorur_arxiv24]. Secondly, all the base scores in Sec. V are set to 0.5, except in the sensitivity analysis. In some scenarios, it may seem reasonable for the system to assign different base scores to users’ arguments. For instance, users might give arguments that are considered very relevant and important to them [Battaglia_icsum24]. Those arguments could have a greater base score. Another factor considered when setting the base scores can be the expertise of the users in the context, e.g. in assistive scenarios, health-related arguments given by the caregiver may be more grounded than those given by the care recipient. While this can improve decision quality, it also raises a risk of systemic bias, prioritising caregiver preferences at the expense of the care recipient’s autonomy. Thirdly, while this study assumes distinct arguments for each user, scenarios where users share arguments but assign different base scores require further exploration. Methods to aggregate individual AFs and base scores into a general AF should be further explored [Dickie_AAMAS25].
Respecting the system’s interaction limitations, we mention the following: firstly, no explanation is generated regarding the reason for the robot’s selection. Explanations could be based on the influence of each argument or their relationships to the final decision (see Sec. V-D). Similarly, if a user complains or questions the reason behind the robot’s decision, the robot must provide a clear and understandable explanation to facilitate the necessary feedback for adjusting its decision. To achieve this, it is essential to identify the necessary and sufficient changes and explanations required within the framework  [Kampik_IJAR24]. Secondly, this work is presented in a particular assistive use case. Even though it is generalisable for any scenario where multiple users have competing preferences, it should be tested and studied in different use cases and in a real-world scenario. Additionally, this work is based on a single-interaction approach, in which a care recipient interacts with the robot once, only to perform the tests. More investigation on how this system behaves in long-term interactions is needed.